OpenStudy (anonymous):

During a paint sale, a hardware store sold 33 gallons of flat paint and 57 gallons of high-gloss paint. What is the ratio of gallons of high-gloss sold to the total gallons sold? A. 57 : 33 B. 33 : 90 C. 57 : 90 D. 33 : 57

OpenStudy (anonymous):

\[\frac{57}{57+33}=\frac{57}{90} \]